The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 37 October 10, 1999

Contents

Abstracts of refereed papers 1
Conference contributions  7
Job opportunities  8
Conference announcement 11

Abstracts of Refereed Papers

Mass-loss rates and luminosity functions of dust-enshrouded AGB stars and red supergiants in the LMC

Jacco Th. van Loon^{1}, M.A.T. Groenewegen^{3}, A. de Koter^{2}, Norman R. Trams^{4}, L.B.F.M. Waters^{2,5}, Albert A. Zijlstra^{6}, Patricia A. Whitelock^{7}, Cecile Loup^{8}

1 Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom 2 Astronomical Institute, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 3 Max-Planck Institut für Astrophysik, Karl-Schwarzschild Straße 1, D-85740 Garching bei München, Germany 4 Astrophysics Division of ESA, ESTEC, P.O.Box 299, NL-2200 AG Noordwijk, The Netherlands 5 Space Research Organization Netherlands, Landleven 12, NL-9700 AV Groningen, The Netherlands 6 University of Manchester Institute of Science and Technology, P.O.Box 88, Manchester M60 1QD, United Kingdom 7 South African Astronomical Observatory, P.O.Box 9, 7935 Observatory, Republic of South Africa 8 Institut d’Astrophysique de Paris, 98bis Boulevard Arago, F-75014 Paris, France

A radiative transfer code is used to model the spectral energy distributions of 57 mass-losing Asymptotic Giant Branch (AGB) stars and red supergiants (RSGs) in the Large Magellanic Cloud (LMC) for which ISO spectroscopic and photometric data are available. As a result we derive mass-loss rates and bolometric luminosities.

A gap in the luminosity distribution around M_{bol} = −7.5 mag separates AGB stars from RSGs. The luminosity distributions of optically bright carbon stars, dust-enshrouded carbon stars and dustenshrouded M-type stars have only little overlap, suggesting that the dust-enshrouded AGB stars are at the very tip of the AGB and will not evolve significantly in luminosity before mass loss ends their AGB evolution.

Derived mass-loss rates span a range from Ṁ ∼ 10^{−7} to 10^{−3} M_{⊙} yr^{−1}. More luminous and cooler stars are found to reach higher mass-loss rates. The highest mass-loss rates exceed the classical limit set by the momentum of the stellar radiation field, L/c, by a factor of a few due to multiple scattering of photons in the circumstellar dust envelope. Mass-loss rates are lower than the mass consumption rate by nuclear burning, Ṁ_{nuc}, for most of the RSGs. Two RSGs have Ṁ ≫ Ṁ_{nuc}, however, suggesting that RSGs shed most of their stellar mantles in short phases of intense mass loss. Stars on the thermal pulsing AGB may also experience episodes of intensified mass loss, but their quiescent mass-loss rates are usually already higher than Ṁ_{nuc}.

Accepted by: Astronomy and Astrophysics Main Journal For preprints, contact jacco@ast.cam.ac.uk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9909416

Catalogue of candidate emission-line objects in the Small Magellanic Cloud

Michael T. Murphy^{1} and Michael S. Bessell^{2}

1 School of Physics, The University of New South Wales, Sydney 2052, Australia 2 Research School of Astronomy and Astrophysics, The Australian National University, Private Bag, Weston Creek P.O., Weston Creek A.C.T. 2611, Australia

Hα and [Oiii] narrow band, wide field (7 × 7 degree), CCD images of the Small Magellanic Cloud were compared and a catalogue of candidate planetary nebulae and Hα emission-line stars was compiled. The catalogue contains 131 planetary nebulae candidates, 23 of which are already known to be or are probable planetary nebulae or very low excitation objects. Also, 218 emission-line candidates have been identified with 113 already known. Our catalogue therefore provides a useful supplement to those of Meyssonnier & Azzopardi (1993) and Sanduleak, MacConnell & Davis Phillip (1978). Further observations are required to confirm the identity of the unknown objects.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact mim@bat.phys.unsw.edu.au
Also available from the URL http://bat.phys.unsw.edu.au/∼mim
or by anonymous ftp at bat.phys.unsw.edu.au , pub/mim/pnpaper.tar.gz

The Large Magellanic Cloud globular cluster NGC 1866: new data, new models, new analysis

V. Testa^{1}, F.R. Ferraro^{2,3}, A. Chieffi^{4}, O. Straniero^{5}, M. Limongi^{1} and F. Fusi Pecci^{6,3}

1 Osservatorio Astronomico di Roma,Via Frascati, 33 00040 Monteporzio Catone, Italy 2 E.S.O., Karl Schwarzschild Strasse 2,D-85748 Garching bei München, Germany 3 Osservatorio Astronomico di Bologna, via Ranzani 1, I-40126 Bologna, Italy 4 Istituto di Astrofisica Spaziale, C.N.R., Via del Fosso del Cavaliere, 00100 Roma, Italy 5 Osservatorio Astronomico di Collurania, Via Maggini, 88 64100 Teramo, Italy 6 Stazione Astronomica di Cagliari, 09012 Capoterra, Italy

We present a new deep (down to V ∼ 24) photometry of a wide region (∼ 6′ × 6′) around the LMC globular cluster NGC1866: our sample is much larger (by more than a factor three) than any previous photometry and with a main sequence which may be considered complete, down to at least 3 mag below the brightest MS star: such an occurrence allows a meaningful and robust comparison with various theoretical scenarios produced by means of models computed with the evolutionary code FRANEC. Both age and present mass function slope, α, are derived by a fit to the available MS and by the use of the parameter Δσ, which is simply the difference, in σ’s, between the observed and predicted integrated MS luminosity functions. Our main conclusions are:

a) the adoption of standard models (i.e. computed by adopting the Schwarzschild criterion to fix the border of the convective core) allows a fair fit to the MS for an age of the order of 100-140 Myr and a present mass function having a slope α between 2.3 and 1.9, the exact values depending on the adopted distance modulus. It is moreover possible to reproduce the average He clump luminosity while the total number of stars predicted in the He clump is twice the observed value: this means that we re-obtain and confirm the first finding of Becker and Mathews (1983), according to whom the simple adoption of a ”classical” scenario leads to a neat discrepancy concerning the prediction of the number of stars in the He clump.

b) the adoption of models computed by increasing the size of the convective core by a certain amount, i.e. 0.25 H_{p}, leads to a fair fit to the main sequence only for a visual distance modulus (m − M)_{V} ≃ 18.6, an age t ≃ 200 Myr and α ≃ 2.2. In this case, the total number of He clump stars is well reproduced, although the luminosity function of the He clump itself is predicted to be systematically less luminous than observed.

The previous conclusions are based on the assumption that there is no appreciable population of binaries in NGC 1866. Though there are not yet sufficient data on the frequency of binary systems in these clusters, we analyzed how the previous scenarios would change if a consistent (≃ 30%) population of binary systems were present in the cluster. This choice is based on the fact that a fraction of binaries of the order of 30% has already been found in NGC 1818, a cluster similar to NGC 1866 (Elson et al., 1998). The inclusion of a 30% binary population leads to the following conclusions:

c) the adoption of the standard models now leads to a good fit to the entire luminosity function, i.e. main sequence, turn off, and He clump stars, for a visual distance modulus (m − M)_{V} = 18.8, an age t ≃ 100 Myr and a mass function slope α ≃ 2.4, thus largely removing the “classical” discrepancy between observed and predicted number of stars in the He burning clump. The quoted visual distance modulus constrains the unreddened distance modulus (m − M)_{0} within 18.50 and 18.62, depending on the reddening (whose most common values available in the literature range from 0.05 to 0.10).

d) at variance with point c), the fit obtained by using models computed with an enlarged convective core gets worse when a binary component is taken into account. This is due to the fact that the presence of binary systems increases the existing discrepancy between the observed and predicted clump luminosity, since the He clump is predicted to be even less luminous than in absence of binaries.

Accepted by: The Astronomical Journal

For preprints, contact testa@coma.mporzio.astro.it
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9909417

The Optical Gravitational Lensing Experiment.

Investigating the Influence of Blending on the Cepheid Distance Scale with Cepheids in the Large Magellanic Cloud

K. Z. Stanek^{1}, A. Udalski^{2}

1 Harvard-Smithsonian Center for Astrophysics, USA 2 Warsaw University Observatory, Poland

We investigate the influence of blending on the Cepheid distance scale. Blending leads to systematically low distances to galaxies observed with HST, and therefore to systematically high estimates of the Hubble constant H_{0}. We select a sample of 43 long-period, large-amplitude Cepheids in the Large Magellanic Cloud, from the catalog recently released by the Optical Gravitational Lensing Experiment. We then model the effects of blending, as observed by HST at large distances, by adding the V, I–bands contributions from nearby bright stars. We find that the derived distance modulus would be too short, compared to the true value, by ∼ 0.07 mag at distance of 12.5 Mpc and by ∼ 0.35 mag at distance of 25 Mpc. This has direct and important implications for the Cepheid distances to galaxies observed by the HST Key Project on the Extragalactic Distance Scale and other teams: half of the KP sample is likely to exhibit a blending bias greater than 0.1 mag.

Submitted to: ApJ Letters

For preprints, contact kstanek@cfa.harvard.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9909346

HST Observations of the LMC Field Around SN 1987A:

Distance Determination With Red Clump and Tip of the Red Giant Branch Stars

M. Romaniello^{1}, M. Salaris^{2}, S. Cassisi^{3} and N. Panagia^{4}

1 ESO, Karl-Schwarzschild-Straße 2, D-85748 Garching bei München, Germany 2 Astrophysics Research Institute, Liverpool John Moores University, Twelve Quays House, Egerton Wharf, Birkenhead CH41 1LD, UK 3 Osservatorio Astronomico di Collurania, Via M. Maggini, 64100 Teramo, Italy 4 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We have used HST-WFPC2 multiband observations of a field around SN 1987A in the Large Magellanic Cloud to measure its distance from the Sun. The observations allowed us to carefully determine the interstellar extinction along the line of sight to a large number of stars and to measure the LMC distance by using two stellar distance indicators: the Red Clump and the Tip of the Red Giant Branch. From an application of the Red Clump method we obtain a distance modulus (m − M)_{0,RC}^{LMC}=18.59±0.04±0.08 mag (statistical plus systematic error), in good agreement with the distance derived by using the Tip of the Red Giant Branch stars, namely (m−M)_{0,T RGB}^{LMC}=18.69±0.25±0.06 mag (statistical plus systematic error). Both values agree well with the distance to the SN 1987A as determined from a study of its inner ring fluorescent echo ((m − M)_{SN 1987A}=18.55±0.05 mag, Panagia 1998), thus excluding distance moduli lower than 18.43 to a 99.7% significance level. Differences with respect to previous results obtained using the same distance indicators are discussed.

Accepted by: The Astrophysical Journal For preprints, contact mromanie@eso.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910082

A New Giant Branch Clump Structure in the Large Magellanic Cloud

A. E. Piatti^{1}, D. Geisler^{1}, E. Bica^{2}, J. J. Clariá^{3}, J. F. C. Santos Jr.^{4}, A. Sarajedini^{5}, H. Dottori^{2}

1 Universidad de Concepción, Departamento de Física, Casilla 160-C, Concepción, Chile 2 Departamento de Astronomia, Instituto de Física, UFRGS,C.P. 15051, 91501-970, Porto Alegre, RS, Brazil 3 Observatorio Astronómico de Córdoba, Laprida 854, 5000, Córdoba, Argentina 4 Dep. de Física, ICEx, UFMG, C.P. 702, 30123-970 Belo Horizonte, MG, Brazil 5 Astronomy Department, Van Vleck Observatory, Wesleyan University, Middletown, CT 06459, USA

We present Washington C, T_{1} CCD photometry of 21 fields located in the northern part of the Large Magellanic Cloud (LMC), and spread over a region of more than 2.5^{□^{o}}, approximately 6^{o} from the bar. The surveyed areas were chosen on the basis of their proximity to SL 388 and SL 509, whose fields showed the presence of a secondary giant clump, observationally detected by Bica et al. (1998, AJ, 116, 723). We also observed NGC 2209, located ∼14^{o} away from SL 509. From the collected data we found that most of the observed field CMDs do not show a separate secondary clump, but rather a continuous vertical structure (VS), which is clearly seen for the first time. The VS also appears in the field of NGC 2209. Its position and size are nearly the same throughout the surveyed regions: it lies below the Red Giant Clump (RGC) and extends from the bottom of the RGC to ∼ 0.45 mag fainter, spanning the bluest color range of the RGC. In two fields as well as in the NGC 2209 field, the RGC is slightly tilted, following approximately the reddening vector, while the VS maintains its verticality. We found that the number of stars in the VS box defined by Δ(C − T_{1}) = 1.45–1.55 mag and ΔT_{1} = 18.75–19.15 mag has a strong spatial variation, reaching the highest VS star density just north-east of SL 509. Moreover, the more numerous the VS stars in a field, the larger the number of LMC giants in the same zone. We also found that, in addition to SL 509, two relatively massive star clusters, SL 515 and NGC 2209, separated by more than ten degrees from each other, develop giant clumps with a considerable number of VS stars. This result demonstrates that VS stars belong to the LMC and are most likely the result of some kind of evolutionary process in the LMC, particularly in those LMC regions with a noticeable large giant population. Our results are successfully predicted by the models of Girardi (1999, MNRAS, submitted) in the sense that a large proportion of 1–2 Gyr old stars mixed with older stars, and with metallicities higher than [Fe/H] ≃ −0.7 should result in a fainter and bluer secondary clump near the mass where degenerate core He burning takes place. However, our results apparently suggest that in order to trigger the formation of VS stars, there should be other conditions besides the appropriate age, metallicity, and the necessary red giant star density. Indeed, stars satisfying the requisites mentioned above are commonly found throughout the LMC, but the VS phenomenon is only clearly seen in some isolated regions. Finally, the fact that clump stars have an intrinsic luminosity dispersion further constrains the use of the clump magnitude as a reliable distance indicator.

Accepted by: The Astronomical Journal

For preprints, contact piatti@gemini.cfm.udec.cl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9909475

The Optical Gravitational Lensing Experiment. Red Clump Stars as a Distance Indicator.

A. Udalski

Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present calibration of the mean I-band brightness of red clump stars on metallicity. Red clump stars were proposed to be a very attractive standard candle for distance determination. The calibration is based on 218 nearby red giant stars whose high quality spectra allowed to determine accurate individual metal abundances. High quality parallaxes (σ_{π}/π < 10%) and photometry of these very bright stars come from Hipparcos measurements.

Metallicity of the sample covers a large range: −0.6 < [Fe/H] < +0.2 dex. We find only a weak dependence of the mean I-band magnitude on metallicity (0.15 mag/dex) in good agreement with previous empirical and theoretical work.

What is more important, the range of metallicity of the Hipparcos sample overlaps with the metallicity of field stars in the LMC, thus making it possible to determine the distance to the LMC by direct comparison of brightness of the local Hipparcos red clump giants with that of LMC stars. Photometry of field red clump giants in four low extinction regions of the LMC halo collected during the OGLE-II microlensing survey compared with the Hipparcos red clump stars data yields the final (free from population effects) ”red-clump method” distance modulus to the LMC: (m − M) = 18.23 ± 0.05 mag. So far, this is the most robust and accurate distance determination to the LMC.

We also present empirical relations between the color indices U − B, B − V and V − I and metallicity of nearby red clump stars.

Submitted to: Ap.J. Letters

For preprints, contact udalski@astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910167

Abstracts of Non-Refereed Papers

Light curves and metal abundances of RR Lyrae variables in the bar of the Large Magellanic Cloud

G. Clementini^{1}, A. Bragaglia^{1}, E. Carretta^{2}, L. Di Fabrizio^{1}, and R. G. Gratton^{2}

1Osservatorio Astronomico di Bologna, Via Ranzani 1, I-40127 Bologna, Italy 2Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy

The Large Magellanic Cloud (LMC) is widely considered a corner-stone of the astronomical distance scale, however a difference of 0.2–0.3 mag exists in its distance as predicted by the short and long distance scales. Distances to the LMC from Population II objects are founded on the RR Lyrae variables. We have undertaken an observational campaign devoted to the definition of the average apparent luminosity and to the study of the mass-metallicity relation for RR Lyraes in the bar of the LMC. These are compared with analogous quantities for cluster RR Lyraes. The purpose is to see whether an intrinsic difference in luminosity, possibly due to a difference in mass, might exist between field and cluster RR Lyraes, which could be responsible for the well-known dichotomy between short and long distance scales. Preliminary results are presented on the V and B−V light curves, the average apparent visual magnitude, and the pulsational properties of 102 RR Lyrae in the bar of the LMC, observed at ESO in January 1999. The photometric data are accurately tied to the Johnson photometric system. Comparison is presented with the photometry of RR Lyraes in the bar of the LMC obtained by the MACHO collaboration (Alcock et al. 1996). Our sample includes 9 double-mode RR Lyraes selected from Alcock et al. (1997) for which an estimate of the metal abundance from the ΔS method is presented.

To appear in: Proceedings of the IAU Colloquium 176, Budapest, Hungary, on “The Impact of Large Scale Surveys on Pulsating Star research”, Eds. L. Szabados & D. Kurtz, ASP Conference Series, in press.

For preprints, contact gisella@astbo3.bo.astro.it
Also available by anonymous ftp at boas5.bo.astro.it/pub/gisella/papers/budatalk.ps

EROS2 microlensing search towards the Magellanic Clouds

T. Lasserre^{1} (on behalf of the EROS collaboration)

1 CEA/Saclay, DSM/DAPNIA/SPP, F-91191 Gif-sur-Yvette, France

eros2 is a second generation microlensing experiment operating since mid-1996 at the European Southern Observatory (ESO) at La Silla (Chile). We present the two year analysis from our microlensing search towards the Small Magellanic Cloud (SMC), and report on the intensive observation of the caustic crossing event MACHO-SMC98-1 and the limit derived on the location of the lens. We also give preliminary results from our search towards the Large Magellanic Cloud (LMC); 25 square degrees are being analyzed and two candidates have been found. This allows us to set another limit on the halo mass fraction comprised of compact objects.

To appear in: ”Gravitational Lensing: Recent Progress and Future Goals”, Boston University, July 1999, ed. T.G. Brainerd and C.S. Kochanek

For preprints, contact lasserre@hep.saclay.cea.fr
Also available from the URL http://xxx.lpthe.jussieu.fr/abs/astro-ph/9909505

Job Opportunities

Postdoctoral Research Associate National Optical Astronomy Observatories

NOAO invites applications to work with Drs. T. Armandroff and G. Jacoby reducing and analyzing Hubble Space Telescope imaging data of Local Group dwarf galaxies. Research will include photometric analysis of time series WFPC2 images to identify and measure light curves for variable stars, such as RR Lyraes, anomalous Cepheids, and Population II Cepheids that serve as probes of stellar populations as well as distance indicators.

We require a Ph.D. in Astronomy and familiarity with IRAF, DAOPHOT, and the scientific areas outlined above. This is initially a two-year position supported by HST grants, with possible extension depending on receipt of additional funds. A starting date in the late spring is preferred, although this is flexible. Application should be made by 30 November 1999 and should include a C.V. and a statement of scientific research interests. Applicants should arrange for letters of recommendation from three scientific references. All application materials should be sent to

Sandra Abbey

NOAO Human Resources Manager

National Optical Astronomy Observatories

P.O. Box 26732

Tucson, AZ 85726

USA

AAE/EOE.

Postdoctoral Fellowships at the University of Bochum and of Bonn

The Graduiertenkolleg The Magellanic System, Galaxy Interaction and the Evolution of Dwarf Galaxies, jointly run at the Universities of Bonn and Bochum, funded by the Deutsche Forschungsgemeinschaft, invites applications for a postdoctoral fellowship. The successful applicant will find her/himself in a pleasant scientific environment, closely collaborating with graduate students on related topics. She/he should be interested in galaxy research, in particular in the investigation of dwarf galaxies. Members of the Graduiertenkolleg (GK) currently pursue both theoretical and observational studies in the field of galaxy research. They have regular access to major radio and optical telescopes around the world and also utilize satellite observatories. The major goal of this graduate or research school is to furnish comprehensive scientific experience and education of graduate students in astrophysics, while working in a team. The scientific focus of the GK is on the star formation history and evolution, on the interstellar medium, and on the structure and kinematics of the Magellanic Clouds as well as of more distant dwarf galaxies (see http://www.astro.uni-bonn.de/~webgk).

Bochum: Preference for the position at Bochum will be given to candidates with a strong background in mm/submm-astronomy and/or radiointerferometry. Information about the fellowship can be obtained from Profs. Ralf-Jürgen Dettmar and Rolf Chini at the Astronomical Institute, Ruhr University in Bochum, Postbox 102148, D-44780 Bochum:

dettmar, chini@astro.ruhr-uni-bochum.de, Tel.: *[49]-234-322-3459 or 322-5802)

Bonn: Preference for the position at Bonn will be given to candidates with a strong background in stellar astronomy. Information about the fellowship can be obtained from Prof. K. S. de Boer, Sternwarte, Univ. Bonn, Tel. *49-228-733656 (address see below).

The DFG is currently funding the GK for the period 1999–2001. The appointment is for one year with a most likely extension for another year. An appointment for a shorter, work-visit like period is also possible. The monthly payment, set by the rules of the DFG, is DM 2,490 to 2,690 (depending on age), plus DM 200 for expenditures, plus a DM 400 allowance for married persons.

Information about the Fellowships can be obtained from the Spokesman (see below).

Applications consisting of CV, list of publications, research plans (max. 1 page) and name of 2 references should arrive as soon as possible (e-mail is fine; send in all cases the announcement of an application by e-mail), and are to be sent to the Spokesman of the Graduiertenkolleg:

Prof. Dr. Klaas S. de Boer

Sternwarte der Universität Bonn

Auf dem Hügel 71

D-53121 Bonn

Germany

   E-Mail: deboer@astro.uni-bonn.de

Doctoral Research Stipendia Magellanic Clouds and Dwarf Galaxies at the Universities of Bochum and Bonn

In the Graduiertenkolleg The Magellanic System, Galaxy Interaction and the Evolution of Dwarf Galaxies, jointly run at the Universities of Bonn and Bochum and funded by the Deutsche Forschungsgemeinschaft, stipendia for doctoral studies are available as of Fall 1999. Research topics are stellar populations, interstellar clouds, dust, and star formation in the MCs and DGs, as well as the kinematical behaviour of such galaxies (see http://www.astro.uni-bonn.de/~webgk).

Information on the Graduate School can be obtained from the Spokesman, Prof. K.S. de Boer in Bonn (tel.: *[49]-228-733656, deboer@astro.uni-bonn.de). Applications are to be sent as soon as possible to:

Prof. K.S. de Boer Spokesman of the GK Sternwarte Auf dem Hügel 71 D-53121 Bonn Germany

Applications include: C.V., copies of relevant exam certificates, an indication of the preferred research area, and names of 2 references.

Meeting Announcement

The Evolution of Galaxies. I. Observational Clues.

Euroconference to be held in Granada (Spain) May 23 – 27, 2000

    E-mail: euroconf@iaa.es
    URLs: http://www.iaa.es/~euroconf/ (Pre-registration)

            http://www.daec.obspm.fr/ThreeConf.html

Important dates:

As soon as possible Pre-registration
15 January 2000  Deadline for final abstract submission
1 March 2000  Second announcement and final list of participants

Scientific Rationale:

The evolution of galaxies has become an observational fact mainly after recent sky surveys (e.g. the HST deep field survey or the Canada-France redshift survey) have shown that the properties of distant galaxies, formed early in the life of the Universe, differ from the nearby ones. New observational windows at X-rays, ultraviolet, infrared and millimetric wavelengths (ROSAT, IUE, IRAS, ISO, IRAM) have revealed that galaxies contain a wealth of components (very hot gas, atomic hydrogen, molecules, dust, dark matter). However, theoretical modelling did not make progress as fast as the census of the content of galaxies. So far, most models were very empirical or semi-empirical. A real understanding of galaxy evolution requires a proper physical description of the galaxy components as well as of the coupling and feedback between them: stars and gas, cold and hot phases of the interstellar medium, large scale and small scale phenomena.

A significant advance is expected in the near future mostly due to the conjunction of two circumstances. New facilities from ground and space (e.g. VLT, FIRST, XMM, etc) will allow one to explore the most distant Universe. During the last decade, new approaches for galaxy modelling have been designed, in which all the relevant large and small scale processes are taken into account in a coherent way. The rapidly developing computing facilities will permit, for the first time, to provide self-consistent models of galaxy evolution.

This is the first of a series of three Euroconferences on the Evolution of Galaxies. It will address the key observations needed for a decisive progress in the understanding of galaxy evolution. Specialists in various fields of Astronomy, observers and theoreticians are expected to participate.

The second conference in October 2001 in St Denis de la Reunion (France) will review the relevant basic building blocks and small-scale processes in galaxy evolution. The third conference in May 2002 in Kiel (Germany) will be devoted to the overall modelling of galaxy evolution.

Scientific Organizing Committee: G. Stasinska (Coordinator), C. Balkowski, J. Beckman, N. Bergvall, F. Combes, U. Fritze-von Alvensleben, J. Gallagher, M. Kontizas, G. Hensler (Chair, 3^{rd}conf.), A. Maeder, B. Pagel, J. Palous, J. Perez-Mercader, J. Silk, R. Terlevich, M. Tosi, L. Vigroux (Chair, 2^{nd}conf.), J.M. Vilchez (Chair, 1^{st}conf.)

Participation:

According to the rules on Euroconferences, attendance is limited to 100 persons. The SOC will select the participants on the basis of their proposed contribution (or motivation for young researchers). Preference will be given to young researchers intending to participate in all the three conferences of the series. Participation of young researchers from European countries is especially encouraged. The SOC will select for oral presentations the contributions dealing explicitly with the topic of this first conference. Posters are welcome, including for the topics related with the themes of the other two conferences of the series. The conference fee will be 200 Euros (1 Euro = 1.07 USD).

Conference Overview and Invited Speakers: Introductory reviews:

The best observables from the point of view of a model maker (Francesca Matteucci) Challenging observations with the new generation of large telescopes (Danielle Alloin)

Element abundances in galaxies (Session 1)

Abundance determinations in stars Abundance determinations in nebulae Abundances distribution in the Milky Way Abundances in spiral, dwarf and starburst galaxies Abundances in elliptical galaxies Keynote lecturers: Monique Spite*, Manuel Peimbert Discussion leaders: Michael Edmunds + Jean-Rene Roy

The interstellar and intergalactic medium (Session 2)

The cold interstellar medium: structure of molecular clouds The warm/hot gaseous component Gas, dust and their relation to star formation in galaxies Keynote lecturers: Francoise Combes, Jurgen Stutzki Discussion leaders: Rolf Chini + Deidre Hunter

Evolution of stellar populations (Session 3)

Analysis of stellar populations by synthesis techniques Disentangling age and metallicity in stellar populations Distribution of stellar masses Keynote lecturers: Gustavo Bruzual, Monica Tosi Discussion leaders: Eduardo Hardy + Angeles Diaz

Structure, dynamics and environment of galaxies (Session 4)

Disk galaxies: structure, bars, assymetries, outflow/inflow Dwarf galaxies: dark matter content, galactic winds, star formation histories, environmental effects Elliptical galaxies: fundamental plane, entropy, clues for formation of ellipticals, ellipticals in

clusters Starburst galaxies: violent star formation, super star clusters Keynote lecturers: Evan Skillman, Eva Grebel, I.F.Mirabel Discussion leaders: Jay Gallagher + R. Terlevich

Galaxies at high redshift (Session 5)

Deep surveys of galaxies Cosmic history of star formation Metallicity indicators at large redshifts Quasars and absorption line systems Keynote lecturers: Rafael Guzman, Max Pettini Discussion leaders: Daniela Calzetti + Mauro Giavalisco

Closing session Perspectives of advancement for the next decade Discussion leader: John Huchra (speakers not yet confirmed are marked with a * )